### v2

@@ 3.2 Closure operators
As mentioned before, our strategy to find the D-ERs of $\mathcal{C}$ will consist of scanning over the entire power-set $2^{\mathcal{E}}$ while ignoring most subsets $\mathcal{X}\subseteq\mathcal{E}$ that are either not down-sets or do not correspond to faces.343434 In this subsection, we will ignore the symmetries and the excluded region $\widehat{\mathcal{C}}$; we will then combine all these ingredients in the following subsection. Since the conditions for $\mathcal{X}$ to be a down-set or a face correspond to two completely different and unrelated structures, namely the cone and the poset, we need a sufficiently general tool that “combines” both of these conditions and make the scan over $2^{\mathcal{E}}$ more efficient. One such tool that will be key in our algorithm is the notion of a closure operator Davey_Priestley_2002 .
@@
which proves the second condition. However, $(\text{cl}_{1}\circ\text{cl}_{2})$ is in general not a closure operator because it does not satisfy the third condition in 3, as $\text{cl}_{2}(\mathcal{X})$ is not necessarily a closed set with respect to $\text{cl}_{1}$.353535 As a simple example, consider a poset $\mathcal{P}$ such that there is no element which is incomparable with every other element. Let $\text{cl}_{1}=\text{cl}_{\text{D}}$ as defined above, and $\text{cl}_{2}=\text{cl}_{\text{U}}$, where $\text{cl}_{\text{U}}$ is the up-set closure operator defined analogously to $\text{cl}_{\text{D}}$ where we simply replace $\preceq$ with $\succeq$ in (31). Then the only subset $\mathcal{X}\subseteq\mathcal{P}$ which is closed under both $\text{cl}_{1}$ and $\text{cl}_{2}$ is $\mathcal{X}=\mathcal{P}$.
@@ 3.3 The algorithm in detail
The computation will be organized in terms of certain D-faces of $\mathcal{C}$ which will be found by intermediate steps of the algorithm. However, since our goal is to compute just the D-ERs, each belonging to multiple faces, only some D-faces will be determined by the algorithm (which D-faces in particular will depend on different choices made during the computation, devised such that the chosen collection contains all the D-ERs). Each D-face $\mathscr{F}$ will be associated to a triplet of the form $\mathfrak{T}=(\mathcal{D},\mathcal{U},G_{\mathcal{D}})$, where $\mathcal{D}=\mathcal{V}(\mathscr{F})$, $G_{\mathcal{D}}$ is the subgroup of $G$ that stabilizes $\mathcal{D}$,363636 This is the stabilizer group of $\mathcal{D}$, i.e., the subgroup of $G$ of elements that leave $\mathcal{D}$ unchanged. and $\mathcal{U}\subseteq\mathcal{E}$ is an up-set specifying an excluded region $\widehat{\mathscr{F}}$ on the boundary of $\mathscr{F}$ as follows. To each ${\sf{E}}_{i}\in\mathcal{U}$, we associate the hyperplane
@@
for some choice of $\sigma_{j}\in\mathbb{R}^{+}$. Since this is a positive linear combination, the inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ is redundant, and it is saturated only by vectors that saturate all inequalities ${\sf{E}}_{j}(\vec{v})\geq 0$ for ${\sf{E}}_{j}\in\mathcal{V}(\mathscr{F}_{i})$. Equivalently, this means that if we let $\mathbb{H}_{i}$ be the hyperplane specified by ${\sf{E}}_{i}$ as in (40), we have $\mathscr{F}_{i}=\mathbb{H}_{i}\cap\mathcal{C}$. This is convenient because it allows us to treat the excluded region $\widehat{\mathcal{C}}$ of $\mathcal{C}$ in the same way as we treat the excluded region $\widehat{\mathscr{F}}$ of a D-face $\mathscr{F}$ obtained at an intermediate stage of the computation. In terms of the poset structure, there are many ways to extend the poset $\mathcal{E}$ to include these new dual vectors ${\sf{E}}_{i}$ (one for each $\mathscr{F}_{i}\in\widehat{\mathcal{C}}$). The simplest is to just have the redundant ${\sf{E}}_{i}$ form an antichain that is “wholly disjoint” from the rest of the poset,373737 By this we mean that each ${\sf{E}}_{i}$ is incomparable to any other element in the poset. although it is oftentimes natural to add the redundant ${\sf{E}}_{i}$ “at the top” of the original poset.383838 By this we mean that each ${\sf{E}}_{i}$ covers some other element of the poset but is not covered by any element. In this case, a simple solution that is guaranteed to preserve the symmetry of the problem is to specify cover relations such that each new ${\sf{E}}_{i}$ covers every maximal element of the original poset. However, this is not strictly necessary, and in specific situations other extensions might be more natural.393939 For example, in the case of the MI-poset, the only redundant inequalities are precisely the ones given by the entropies “at the top” of the poset, and they specify part of the excluded region. Importantly, regardless of how we choose to extend the poset with the new dual vectors, while the full set of D-ERs of $\mathcal{C}$ might change because of the extension, the set of D-ERs which are not in $\widehat{\mathcal{C}}$ remains unchanged.
@@
Typically, $f$ will be a function that can be computed efficiently from the data given by a triplet, and should be chosen depending on the specific details of the problem to which the algorithm is applied. One useful example of this function is an upper bound to the dimension of the face corresponding to the triplet $\mathfrak{T}$, where $f$ returns TRUE if the bound is satisfied and FALSE otherwise. If one wants to use the algorithm to directly find all $G$-orbits of D-ERs of $\mathcal{C}$ which are not in $\widehat{\mathcal{C}}$, then one can simply set this bound to 1. Another option instead is to set a larger upper bound for the dimension of a face, and use the algorithm to restrict the search space to a point where one can then use other methods to find the extreme rays of the resulting lower-dimensional faces. An alternative option that might be useful is to stop the computation only when the excluded region on a face is sufficiently vast.404040 We will use the first option to obtain the results presented in §4 and a combination of the latter two to obtain the results presented in §5.
@@
We now describe the conditions that we require for each triplet and the corresponding face. By our assumptions about the set-up described in §3.1, it is trivial to check that the triplet $\mathfrak{T}^{(0)}$ used to initialize the algorithm satisfies these conditions.414141 This is true except for trivial situations, such as the case where $\mathcal{C}$ is 1-dimensional, which we ignore.
@@
    After constructing $\overline{\mathcal{U}}$, one may wonder if the efficiency of the algorithm can be improved by further checking whether there is some $\mathcal{D}_{m}\in\dot{\mathfrak{D}}$ such that $\mathcal{D}_{m}\cap\overline{\mathcal{U}}\neq\varnothing$ while $\mathcal{D}_{m}\cap\mathcal{U}=\varnothing$, in which case one could add more elements to $\overline{\mathcal{U}}$ and extend the excluded region for the new triplets that will be constructed in the following steps. However, this is not the case due to the following implication: 424242 We can prove this as follows. Let $\mathcal{D}_{m}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m}\})$ for some ${\sf{E}}_{m}\in\mathcal{M}_{m}$, and suppose $\mathcal{D}_{m}$ contains an element $\sf{U}\in\mathcal{M}_{m^{\prime}}$ for some $\mathcal{M}_{m^{\prime}}\subseteq\overline{\mathcal{U}}\setminus\mathcal{U}$. From the algorithm, this means that we have constructed some $\mathcal{D}_{m^{\prime}}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m^{\prime}}\})$ with ${\sf{E}}_{m^{\prime}}\in\mathcal{M}_{m^{\prime}}$ such that there is some $\sf{U}^{\prime}\in\mathcal{D}_{m^{\prime}}\cap\mathcal{U}$. By the defining properties of a closure operator, $\mathcal{D}_{m}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m},\sf{U}\})\supseteq\text{cl}_{\text{FD}}(\mathcal{D}\cup\{\sf{U}\})$, and since $\mathcal{M}_{m^{\prime}}$ is an orbit of $\mathcal{M}$ generated by $G_{\mathcal{D}}$, there is some $g\in G_{\mathcal{D}}$ such that $g(\sf{U})={\sf{E}}_{m^{\prime}}$. To show that $\mathcal{D}_{m}\cap\mathcal{U}\neq\varnothing$, it is sufficient to show that $g(\mathcal{D}_{m}\cap\mathcal{U})\neq\varnothing$. Since $\mathcal{U}$ is invariant under $G_{\mathcal{D}}$, we have $g(\mathcal{D}_{m}\cap\mathcal{U})=g(\mathcal{D}_{m})\cap\mathcal{U}$, and by the above inclusion, $g(\mathcal{D}_{m})\cap\mathcal{U}\supseteq g(\text{cl}_{\text{FD}}(\mathcal{D}\cup\{\sf{U}\}))\cap\mathcal{U}$. Furthermore, since $g$ and $\text{cl}_{\text{FD}}$ commute (cf. (39)), and $\mathcal{D}$ is invariant under $g$, we obtain $g(\mathcal{D}_{m}\cap\mathcal{U})\supseteq\text{cl}_{\text{FD}}(\mathcal{D}\cup g(\{\sf{U}\}))\cap\mathcal{U}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m^{\prime}}\}\cap\mathcal{U}=\mathcal{D}_{m^{\prime}}\cap\mathcal{U}\ni\sf{U}^{\prime}$.
@@
Consider an element ${\sf{E}}\in\mathcal{M}$ and the set $\mathcal{D}\cup\{{\sf{E}}\}$. If there is another element ${\sf{E}}^{\prime}$ which is mapped to ${\sf{E}}$ by an element of the stabilizer group $G_{\mathcal{D}}$ of $\mathcal{D}$, then clearly $\mathcal{D}\cup\{{\sf{E}}\}$ and $\mathcal{D}\cup\{{\sf{E}}^{\prime}\}$ are related by the same group operation.434343 Recall that the action of $G$ on subsets of $\mathcal{E}$ is defined element-wise, starting from that on $\mathcal{E}$. Therefore, since we are only interested in finding D-ERs up to equivalence under the action of $G_{\mathcal{D}}$, once we include in the scan the down-sets obtained from $\mathcal{D}$ by adding ${\sf{E}}$, we can ignore any down-set obtained from $\mathcal{D}$ by adding ${\sf{E}}^{\prime}$. We then partition $\mathcal{M}$ into $G_{\mathcal{D}}$-orbits and focus only on deformations of $\mathcal{D}$ generated by representatives of these orbits. This partition of $\mathcal{M}$ is denoted by $\mathfrak{M}$.
@@
Nevertheless, we stress that even when using standard conversion algorithms, it might still be possible to speed up this computation significantly using the set $\mathcal{U}$ that specifies the excluded region.444444 This trick was indeed crucial for the derivation of the results presented in §5. To see this, consider a triplet $(\mathcal{D},\mathcal{U},G_{\mathcal{D}})$ for a face $\mathscr{F}$, and suppose that $\mathcal{U}$ is a “large subset” of $\mathcal{E}$, i.e., that $|\mathcal{U}|/|\mathcal{E}|\sim 1$. On the subspace spanned by $\mathscr{F}$, which is determined by $\mathcal{D}$, $\mathscr{F}$ is a pointed cone carved out by the inequalities corresponding to the elements of $\mathcal{E}\setminus\mathcal{D}$. On this subspace, consider the larger cone $\check{\mathscr{F}}$ carved out by the inequalities given by $\mathcal{E}\setminus(\mathcal{D}\cup\mathcal{U})$. Notice that since the triplet we are considering is in the output of the algorithm, it satisfies in particular (iv) in Condition 1, implying that this cone is pointed. While the ERs of $\check{\mathscr{F}}$ are in general different from those of $\mathscr{F}$, notice that any ER of $\mathscr{F}$ which does not saturate any inequality from $\mathcal{U}$ (i.e., that is not in the excluded region $\widehat{\mathscr{F}}$) is also an ER of $\check{\mathscr{F}}$. Since $\check{\mathscr{F}}$ is specified by a much smaller set of inequalities compared to $\mathscr{F}$, finding its ERs using standard conversion algorithm can be much simpler.
@@ 3.4 A few possible variations
In some situations one might be interested in finding only the D-ERs that belong to a lower-dimensional face $\mathscr{F}$ of the cone. For example, this is the case for the SAC, where as a result of 1, we can focus on the face $\mathscr{F}_{*}$ that we mentioned in the Introduction. In these situations, it would seem natural to first simplify the problem via a dimensional reduction to the subspace spanned by $\mathscr{F}$, and then apply the algorithm to the resulting set-up. However, as we discussed above when we commented on the post-processing of the ERs, the full structure of the poset is not “completely visible” in a lower-dimensional subspace. One practical alternative is to instead initialize the algorithm setting $\mathcal{D}^{(0)}=\mathcal{V}(\mathscr{F})$ instead of $\mathcal{D}^{(0)}=\varnothing$.454545 Here we are assuming that $\mathscr{F}$ is a D-face. If it is not, then one should simply set $\mathcal{D}^{(0)}=\text{cl}_{\text{FD}}(\mathcal{V}(\mathscr{F}))$.

### v3

@@ 3.2 Closure operators
As mentioned before, our strategy to find the D-ERs of $\mathcal{C}$ will consist of scanning over the entire power-set $2^{\mathcal{E}}$ while ignoring most subsets $\mathcal{X}\subseteq\mathcal{E}$ that are either not down-sets or do not correspond to faces.333333 In this subsection, we will ignore the symmetries and the excluded region $\widehat{\mathcal{C}}$; we will then combine all these ingredients in the following subsection. Since the conditions for $\mathcal{X}$ to be a down-set or a face correspond to two completely different and unrelated structures, namely the cone and the poset, we need a sufficiently general tool that “combines” both of these conditions and make the scan over $2^{\mathcal{E}}$ more efficient. One such tool that will be key in our algorithm is the notion of a closure operator Davey_Priestley_2002 .
@@
which proves the second condition. However, $(\text{cl}_{1}\circ\text{cl}_{2})$ is in general not a closure operator because it does not satisfy the third condition in 3, as $\text{cl}_{2}(\mathcal{X})$ is not necessarily a closed set with respect to $\text{cl}_{1}$.343434 As a simple example, consider a poset $\mathcal{P}$ such that there is no element which is incomparable with every other element. Let $\text{cl}_{1}=\text{cl}_{\text{D}}$ as defined above, and $\text{cl}_{2}=\text{cl}_{\text{U}}$, where $\text{cl}_{\text{U}}$ is the up-set closure operator defined analogously to $\text{cl}_{\text{D}}$ where we simply replace $\preceq$ with $\succeq$ in (31). Then the only subset $\mathcal{X}\subseteq\mathcal{P}$ which is closed under both $\text{cl}_{1}$ and $\text{cl}_{2}$ is $\mathcal{X}=\mathcal{P}$.
@@ 3.3 The algorithm in detail
The computation will be organized in terms of certain D-faces of $\mathcal{C}$ which will be found by intermediate steps of the algorithm. However, since our goal is to compute just the D-ERs, each belonging to multiple faces, only some D-faces will be determined by the algorithm (which D-faces in particular will depend on different choices made during the computation, devised such that the chosen collection contains all the D-ERs). Each D-face $\mathscr{F}$ will be associated to a triplet of the form $\mathfrak{T}=(\mathcal{D},\mathcal{U},G_{\mathcal{D}})$, where $\mathcal{D}=\mathcal{V}(\mathscr{F})$, $G_{\mathcal{D}}$ is the subgroup of $G$ that stabilizes $\mathcal{D}$,353535 This is the stabilizer group of $\mathcal{D}$, i.e., the subgroup of $G$ of elements that leave $\mathcal{D}$ unchanged. and $\mathcal{U}\subseteq\mathcal{E}$ is an up-set specifying an excluded region $\widehat{\mathscr{F}}$ on the boundary of $\mathscr{F}$ as follows. To each ${\sf{E}}_{i}\in\mathcal{U}$, we associate the hyperplane
@@
for some choice of $\sigma_{j}\in\mathbb{R}^{+}$. Since this is a positive linear combination, the inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ is redundant, and it is saturated only by vectors that saturate all inequalities ${\sf{E}}_{j}(\vec{v})\geq 0$ for ${\sf{E}}_{j}\in\mathcal{V}(\mathscr{F}_{i})$. Equivalently, this means that if we let $\mathbb{H}_{i}$ be the hyperplane specified by ${\sf{E}}_{i}$ as in (40), we have $\mathscr{F}_{i}=\mathbb{H}_{i}\cap\mathcal{C}$. This is convenient because it allows us to treat the excluded region $\widehat{\mathcal{C}}$ of $\mathcal{C}$ in the same way as we treat the excluded region $\widehat{\mathscr{F}}$ of a D-face $\mathscr{F}$ obtained at an intermediate stage of the computation. In terms of the poset structure, there are many ways to extend the poset $\mathcal{E}$ to include these new dual vectors ${\sf{E}}_{i}$ (one for each $\mathscr{F}_{i}\in\widehat{\mathcal{C}}$). The simplest is to just have the redundant ${\sf{E}}_{i}$ form an antichain that is “wholly disjoint” from the rest of the poset,363636 By this we mean that each ${\sf{E}}_{i}$ is incomparable to any other element in the poset. although it is oftentimes natural to add the redundant ${\sf{E}}_{i}$ “at the top” of the original poset.373737 By this we mean that each ${\sf{E}}_{i}$ covers some other element of the poset but is not covered by any element. In this case, a simple solution that is guaranteed to preserve the symmetry of the problem is to specify cover relations such that each new ${\sf{E}}_{i}$ covers every maximal element of the original poset. However, this is not strictly necessary, and in specific situations other extensions might be more natural.383838 For example, in the case of the MI-poset, the only redundant inequalities are precisely the ones given by the entropies “at the top” of the poset, and they specify part of the excluded region. Importantly, regardless of how we choose to extend the poset with the new dual vectors, while the full set of D-ERs of $\mathcal{C}$ might change because of the extension, the set of D-ERs which are not in $\widehat{\mathcal{C}}$ remains unchanged.
@@
Typically, $f$ will be a function that can be computed efficiently from the data given by a triplet, and should be chosen depending on the specific details of the problem to which the algorithm is applied. One useful example of this function is an upper bound to the dimension of the face corresponding to the triplet $\mathfrak{T}$, where $f$ returns TRUE if the bound is satisfied and FALSE otherwise. If one wants to use the algorithm to directly find all $G$-orbits of D-ERs of $\mathcal{C}$ which are not in $\widehat{\mathcal{C}}$, then one can simply set this bound to 1. Another option instead is to set a larger upper bound for the dimension of a face, and use the algorithm to restrict the search space to a point where one can then use other methods to find the extreme rays of the resulting lower-dimensional faces. An alternative option that might be useful is to stop the computation only when the excluded region on a face is sufficiently vast.393939 We will use the first option to obtain the results presented in §4 and a combination of the latter two to obtain the results presented in §5.
@@
We now describe the conditions that we require for each triplet and the corresponding face. By our assumptions about the set-up described in §3.1, it is trivial to check that the triplet $\mathfrak{T}^{(0)}$ used to initialize the algorithm satisfies these conditions.404040 This is true except for trivial situations, such as the case where $\mathcal{C}$ is 1-dimensional, which we ignore.
@@
    After constructing $\overline{\mathcal{U}}$, one may wonder if the efficiency of the algorithm can be improved by further checking whether there is some $\mathcal{D}_{m}\in\dot{\mathfrak{D}}$ such that $\mathcal{D}_{m}\cap\overline{\mathcal{U}}\neq\varnothing$ while $\mathcal{D}_{m}\cap\mathcal{U}=\varnothing$, in which case one could add more elements to $\overline{\mathcal{U}}$ and extend the excluded region for the new triplets that will be constructed in the following steps. However, this is not the case due to the following implication: 414141 We can prove this as follows. Let $\mathcal{D}_{m}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m}\})$ for some ${\sf{E}}_{m}\in\mathcal{M}_{m}$, and suppose $\mathcal{D}_{m}$ contains an element $\sf{U}\in\mathcal{M}_{m^{\prime}}$ for some $\mathcal{M}_{m^{\prime}}\subseteq\overline{\mathcal{U}}\setminus\mathcal{U}$. From the algorithm, this means that we have constructed some $\mathcal{D}_{m^{\prime}}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m^{\prime}}\})$ with ${\sf{E}}_{m^{\prime}}\in\mathcal{M}_{m^{\prime}}$ such that there is some $\sf{U}^{\prime}\in\mathcal{D}_{m^{\prime}}\cap\mathcal{U}$. By the defining properties of a closure operator, $\mathcal{D}_{m}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m},\sf{U}\})\supseteq\text{cl}_{\text{FD}}(\mathcal{D}\cup\{\sf{U}\})$, and since $\mathcal{M}_{m^{\prime}}$ is an orbit of $\mathcal{M}$ generated by $G_{\mathcal{D}}$, there is some $g\in G_{\mathcal{D}}$ such that $g(\sf{U})={\sf{E}}_{m^{\prime}}$. To show that $\mathcal{D}_{m}\cap\mathcal{U}\neq\varnothing$, it is sufficient to show that $g(\mathcal{D}_{m}\cap\mathcal{U})\neq\varnothing$. Since $\mathcal{U}$ is invariant under $G_{\mathcal{D}}$, we have $g(\mathcal{D}_{m}\cap\mathcal{U})=g(\mathcal{D}_{m})\cap\mathcal{U}$, and by the above inclusion, $g(\mathcal{D}_{m})\cap\mathcal{U}\supseteq g(\text{cl}_{\text{FD}}(\mathcal{D}\cup\{\sf{U}\}))\cap\mathcal{U}$. Furthermore, since $g$ and $\text{cl}_{\text{FD}}$ commute (cf. (39)), and $\mathcal{D}$ is invariant under $g$, we obtain $g(\mathcal{D}_{m}\cap\mathcal{U})\supseteq\text{cl}_{\text{FD}}(\mathcal{D}\cup g(\{\sf{U}\}))\cap\mathcal{U}=\text{cl}_{\text{FD}}(\mathcal{D}\cup\{{\sf{E}}_{m^{\prime}}\}\cap\mathcal{U}=\mathcal{D}_{m^{\prime}}\cap\mathcal{U}\ni\sf{U}^{\prime}$.
@@
Consider an element ${\sf{E}}\in\mathcal{M}$ and the set $\mathcal{D}\cup\{{\sf{E}}\}$. If there is another element ${\sf{E}}^{\prime}$ which is mapped to ${\sf{E}}$ by an element of the stabilizer group $G_{\mathcal{D}}$ of $\mathcal{D}$, then clearly $\mathcal{D}\cup\{{\sf{E}}\}$ and $\mathcal{D}\cup\{{\sf{E}}^{\prime}\}$ are related by the same group operation.424242 Recall that the action of $G$ on subsets of $\mathcal{E}$ is defined element-wise, starting from that on $\mathcal{E}$. Therefore, since we are only interested in finding D-ERs up to equivalence under the action of $G_{\mathcal{D}}$, once we include in the scan the down-sets obtained from $\mathcal{D}$ by adding ${\sf{E}}$, we can ignore any down-set obtained from $\mathcal{D}$ by adding ${\sf{E}}^{\prime}$. We then partition $\mathcal{M}$ into $G_{\mathcal{D}}$-orbits and focus only on deformations of $\mathcal{D}$ generated by representatives of these orbits. This partition of $\mathcal{M}$ is denoted by $\mathfrak{M}$.
@@
Nevertheless, we stress that even when using standard conversion algorithms, it might still be possible to speed up this computation significantly using the set $\mathcal{U}$ that specifies the excluded region.434343 This trick was indeed crucial for the derivation of the results presented in §5. To see this, consider a triplet $(\mathcal{D},\mathcal{U},G_{\mathcal{D}})$ for a face $\mathscr{F}$, and suppose that $\mathcal{U}$ is a “large subset” of $\mathcal{E}$, i.e., that $|\mathcal{U}|/|\mathcal{E}|\sim 1$. On the subspace spanned by $\mathscr{F}$, which is determined by $\mathcal{D}$, $\mathscr{F}$ is a pointed cone carved out by the inequalities corresponding to the elements of $\mathcal{E}\setminus\mathcal{D}$. On this subspace, consider the larger cone $\check{\mathscr{F}}$ carved out by the inequalities given by $\mathcal{E}\setminus(\mathcal{D}\cup\mathcal{U})$. Notice that since the triplet we are considering is in the output of the algorithm, it satisfies in particular (iv) in Condition 1, implying that this cone is pointed. While the ERs of $\check{\mathscr{F}}$ are in general different from those of $\mathscr{F}$, notice that any ER of $\mathscr{F}$ which does not saturate any inequality from $\mathcal{U}$ (i.e., that is not in the excluded region $\widehat{\mathscr{F}}$) is also an ER of $\check{\mathscr{F}}$. Since $\check{\mathscr{F}}$ is specified by a much smaller set of inequalities compared to $\mathscr{F}$, finding its ERs using standard conversion algorithm can be much simpler.
@@ 3.4 A few possible variations
In some situations one might be interested in finding only the D-ERs that belong to a lower-dimensional face $\mathscr{F}$ of the cone. For example, this is the case for the SAC, where as a result of 1, we can focus on the face $\mathscr{F}_{*}$ that we mentioned in the Introduction. In these situations, it would seem natural to first simplify the problem via a dimensional reduction to the subspace spanned by $\mathscr{F}$, and then apply the algorithm to the resulting set-up. However, as we discussed above when we commented on the post-processing of the ERs, the full structure of the poset is not “completely visible” in a lower-dimensional subspace. One practical alternative is to instead initialize the algorithm setting $\mathcal{D}^{(0)}=\mathcal{V}(\mathscr{F})$ instead of $\mathcal{D}^{(0)}=\varnothing$.444444 Here we are assuming that $\mathscr{F}$ is a D-face. If it is not, then one should simply set $\mathcal{D}^{(0)}=\text{cl}_{\text{FD}}(\mathcal{V}(\mathscr{F}))$.